# F5R-TTS: Improving Flow Matching based Text-to-Speech with Group Relative Policy Optimization

### v1

@@ 4 Conclusion
In this paper, we propose F5R-TTS, which introduces the GRPO method into a flow matching-based NAR TTS system. By transforming the outputs of flow-matching based model into probabilistic representations, GRPO is thus able to be integrated into the training pipeline. Experimental results demonstrate that the proposed method achieves higher SIM and lower WER compared to baseline systems, indicating that GRPO with appropriate reward functions positively contributes to both semantic accuracy and speaker similarity.
@@
  • Reward function optimization: To further enhance model stability in challenging cases, we will continue to investigate optimized reward function designs.
  • Data exploration: In order to delve into how the model performs when exposed to a larger volume of data, we will leverage more training data to conduct further experiments.

### v2

@@ 4 Conclusion
In this paper, we propose F5R-TTS, which introduces the GRPO method into a flow matching-based NAR TTS system. By transforming the outputs of flow-matching based model into probabilistic representations, GRPO can be integrated into the training pipeline. Experimental results demonstrate that the proposed method achieves higher SIM and lower WER compared to baseline systems, indicating that GRPO with appropriate reward functions positively contributes to both semantic accuracy and speaker similarity.
@@
  • Reward function optimization: In order to further enhance the model’s stability in challenging scenarios, we will continue to investigate optimized reward function designs.
  • Data exploration: To better understand the model’s performance with larger datasets, we will utilize more training data for further experiments.